# Controllable Learning in Information Retrieval: A Survey

### v1

@@ 7 Future Directions
Controllable learning is gaining increasing attention, even though much of the current focus remains within implicit domains. This section highlights several future development directions, based on the challenging issues and emerging technologies (e.g., large language models). It is hoped that this chapter will provide valuable inspiration for future research. Why should we control? Theoretical analyses of controllable learning. Finding the optimal hypothesis in the hypothesis space for controllable learning is a more challenging task compared to traditional machine learning. Specifically, given the task requirements, the essence of in-processing controllable learning methods lies in establishing a mapping between the task target and the model parameters. Due to the vast parameter space of current deep learning models, uncovering structural information within this space and understanding causal associations with the target require rigorous theoretical analysis and effective training methods. This stands as a critical future research direction in controllable learning.
Controllable decision-making models. In numerous practical applications such as IR, feedback often takes the form of bandit feedback (i.e., feedback on decisions not executed is unobserved). This presents new challenges for sequential decision-making models such as reinforcement learning. Balancing exploration and exploitation while achieving adaptive control over task requirements is a crucial issue in both the theoretical analysis and practical applications of controllable learning.
@@
Cost-effective control learning mechanisms. As stated in the definition of controllable learning (Definition 1), compared to the original learner, a controllable learner requires additional assistance from control functions, which inevitably introduces additional computational cost. Given the substantial computational costs linked with large-scale models, investigating efficient and cost-effective control mechanisms becomes imperative for future research.
@@ 8 Conclusion
The landscape of controllable learning (CL) has been significantly enriched through the integration of diverse methodologies aimed at enhancing the trustworthiness of machine learning. CL’s ability to dynamically adjust learner according to predefined targets and adapt without retraining when those targets evolve makes it a critical component in fostering reliable and adaptive machine learning models. Particularly within the context of information retrieval (IR), CL offers a means to address the complexity and dynamism inherent in information needs. However, there remains much to explore in terms of theoretical guarantees, computational efficiency, and the integration of LLMs into controllable learning. Further advancements in these areas could lead to more sophisticated, user-centric, and adaptable AI models that meet the diverse and evolving needs of IR applications.

### v2

@@ 5.2 Datasets
In fact, there are currently no dedicated domain-specific datasets for controllable learning. Developing more tailored datasets remains an open challenge, which we leave for future work.
@@ 7 Future Directions
Controllable learning is gaining increasing attention, even though much of the current focus remains within implicit domains. This section highlights several future development directions, based on the challenging issues and emerging technologies (e.g., large language models). It is hoped that this chapter will provide valuable inspiration for future research.
Why should we control? Theoretical analyses of controllable learning. Finding the optimal hypothesis in the hypothesis space for controllable learning is a more challenging task compared to traditional machine learning. Specifically, given the task requirements, the essence of in-processing controllable learning methods lies in establishing a mapping between the task target and the model parameters. Due to the vast parameter space of current deep learning models, uncovering structural information within this space and understanding causal associations with the target require rigorous theoretical analysis and effective training methods. This stands as a critical future research direction in controllable learning.
Controllable sequential decision-making models. In numerous streaming applications such as streaming recommender systems, feedback often takes the form of bandit feedback (i.e., feedback on decisions not executed is unobserved). This presents new challenges for sequential decision-making models such as reinforcement learning and online learning. Balancing exploration and exploitation while achieving adaptive control over task requirements is a crucial issue in both the theoretical analysis and practical applications of controllable learning.
@@
Cost-effective control learning mechanisms. As stated in the definition of controllable learning (Definition 2.1), compared to the original learner, a controllable learner requires additional assistance from control functions, which inevitably introduces additional computational cost. Given the substantial computational costs linked with large-scale models, investigating efficient and cost-effective control mechanisms becomes imperative for future research.
@@ 8 Conclusion
The landscape of controllable learning (CL) has been significantly enriched through the integration of diverse methodologies aimed at enhancing the trustworthiness of machine learning. CL’s ability to dynamically adjust learner according to predefined targets and adapt without retraining when those targets evolve makes it a critical component in fostering reliable and adaptive machine learning models. Particularly within the context of information retrieval (IR), CL offers a means to address the complexity and dynamism inherent in information needs. However, there remains much to explore in terms of theoretical guarantees, computational efficiency, and the integration of LLMs into controllable learning. Further advancements in these areas could lead to more sophisticated, user-centric, and controllable AI models that meet the diverse and evolving needs of IR applications.